Stellar Company offers an MP3 download (seven-single medley) as a premium for every 5 candy bar wrappers presented by customers together with $2.55. The candy bars are sold by the company to distributors for 30 cents each. The purchase price of each download code to the company is $2.30. In addition, it costs 50 cents to distribute each code. The results of the premium plan for the years 2017 and 2018 are as follows. (All purchases and sales are for cash.) 2017 2018 MP3 codes purchased 255,000 336,600 Candy bars sold 3,147,300 2,880,400 Wrappers redeemed 1,224,000 1,530,000 2017 wrappers expected to be redeemed in 2018 295,800 2018 wrappers expected to be redeemed in 2019 357,000 1-Prepare the journal entries that should be made in 2017 and 2018 to record the transactions related to the premium plan of the Stellar Company. (If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually. Round answers to 0 decimal places, e.g. 1,525.) 2-Indicate the amounts for each accounts, and classifications of the items related to the premium plan that would appear on the balance sheet and the income statement at the end of 2017 and 2018.
Answer #1

1.

Account Titles and Explanation Debit Credit
2017
Inventory of premiums (255000 x $2.30) 586500
Cash 586500
(To record the premium inventory)
Cash (3147300 x $0.30) 944190
Sales revenue 944190
(To record the sales)
Cash [1224000/5 = 244800 x ($2.55 - $0.50)] 501840
Premium Expense [244800 x ($2.30 + $0.50 - $2.55)] 61200
Inventory of Premiums (244800 x $2.30) 563040
(To record the expense associated with the sale)
Premium expense 14790
Premium liability [295800/5 = 59160 x ($2.30 + $0.50 - $2.55)] 14790
(To record the premium liability)
2018
Inventory of Premiums (336600 x $2.30) 774180
Cash 774180
(To record the premium inventory)
Cash (2880400 x $0.30) 864120
Sales Revenue 864120
(To record the sales)
Cash [1530000/5 = 306000 x ($2.55 - $0.50)] 627300
Premium Expense [[306000 x ($2.30 + $0.50 - $2.55)] - $14790] 61710
Premium Liability 14790
Inventory of Premiums (306000 x $2.30) 703800
(To record the expense associated with the sale)
Premium Expense 17850
Premium Liability [357000/5 = 71400 x ($2.30 + $0.50 - $2.55)] 17850
(To record the premium liability)

2.

Account Classification Amount
2017 2018
Inventory of Premiums Balance Sheet - Current Asset 23460 93840
Premium Liability Balance Sheet - Current Liability 14790 17850
Premium Expense Income Statement - Selling Expenses 75990 79560